A Complete Bibliography of Publications in Transactions of the American Fisheries

Society: 1920–1929

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

09 June 2023 Version 1.01

Title word cross-reference

10th [Ano20e]. 1920 [Ano21f]. 1921 [Bow22a]. 1922 [Ano23h]. 1923. [Ano23i]. 1924 [Ano24f]. 1925 [Ano25f]. 1926 [Ano26f]. 1927 [Ano27f]. 1928 [Ano28h]. 1929 [Ano29a].

8th [Ano20e].

9th [Ano20e].

Achievement [O’M28]. Acipenser [Bor25]. Active [Ano24a]. Activities [Rad21]. Addition [DJ24]. Adequate [Mil21]. Adjustment [Bro23a]. Administration [Pri21]. Advantage [EH25]. Aeration [Moo29]. Affecting [Mer24]. affinis [BA21]. Age [Bar24, Bor24]. Aid [Pri23]. Alabama [Ano26f]. Alaska [O’M28, Smi21a]. Alewife [Bel20]. Algae [Kid27]. Algal [Til29]. Allentown [Bow22a]. Alosa [Bor24]. ambloplitis [Ban27b, Moo26a]. amended [Ano20b, Ano21b, Ano22b, Ano23b, Ano24c, Ano25b, Ano26b, Ano27b, Ano28c, Ano29c]. American [Ano20e, Ano21f, Ano23h, Ano23i, Ano24f, Ano25f, Ano26f, Ano27f, Ano28h, Ano29a, Bow22a, Bab28]. Analysis [Fie22]. Anatomical [Ken21b]. Angler [Loc29]. Angling [Hen23]. Animal [Nur28, Nur29, TCCM29]. Annual [Ano20e, Ano21f, Ano23h, Ano23i, Ano24f, Ano25f, Ano26f, Ano27f, Ano28h, Ano29a, Bow22a]. Antagonism [Pow21]. Aquarium [Gor26]. Aquariums [Bow22b]. Aquatic [BA21]. Aquiculture [Rad26, Rad27]. argus [Cra21]. Artesian [Hen24]. Artificial [Cra21, Ell29, EG24, Hay27, Mel27, Osb23, Reu20, Wel21]. Ascend [Pri20]. Asset [Smi21a]. Assist [Web29b]. Association [Syn28]. Associations [Nur29]. Atlantic [Bor25, Hun21, LeC26, War21a]. Attending [Sur25]. August [Ano25f, Ano27f, Ano28h].

Back [Tho25]. Bacteriological [Fie22]. Bait [Ban29]. Banquet [Ano28a, Moo28]. Barbed [Doz25]. Barren [Rod24a]. Basic [TCCM29]. Bass [Bal26, Ban27b, Bee24, Bul23, Dav28b, Hay27, Hes29, LeC26, LeC28, Lyd25, Lyd26a, Lyd26b, Moo25, Moo26a, Sco28, Sny23, Sny26, TK21, Wic21]. Basses [Osb23]. Bay [Den20, Mav21]. Beneficial [Hay20]. Benefit [Van25]. Better [Mil21]. between [Loc29]. Biological [Bor25, Dav29, Kei23, McC29, Moo27, Sur23]. Black [Bal26, Bee24, Lyd25, Lyd26b, Osb23, Sny23, Sny26, TK21, Wic21]. Blackhead [Lor27]. Blue [Ken21a, Lyd26a]. Bluegills [Bro26]. Board [But24]. Borodin [Bar24]. Bream [Bro26]. Breeders [EH25, Kei22]. Brief [Coo29, Lyd23]. British [Dym28, Mun28]. Brook [Bin27, Din27, EG24, EH25, MBD27, TCCM28, TCCM29]. Bureau [Rad21, Coo29, Cul23]. Burlington [Bin27]. By-Laws [Ano25b, Ano26b, Ano27b, Ano28c, Ano29c]. By-products [TCCM29].

California [Sco28]. called [Ken21a]. Can [Bab21]. Canada [Ano21f, Ano24f, Bab21, Pri21, Rod24b]. Canadian [Bab28]. Canadian-American [Bab28]. Canals [Pri23]. Canned [Fie22]. Canning [Bea23]. Capacity [BA21]. Capture [Hes29]. Carbohydrates [TCCM29]. Care [Hoo29]. Catch [Tul26]. Catfish [Cla29]. Certain [BA21, Emb22b, Fos25, Kei23, Sur25]. Certificate [Ano20a, Ano21a, Ano22a, Ano23a, Ano24b, Ano25a, Ano26a, Ano27a, Ano28b, Ano29b]. Cestode [Ban27b]. Champaign [Tho29a]. Channel [Cla29]. Chesapeake [Den20]. Chief [Pau24]. Chinook [NL28]. Chironomid [Joh29]. Circulation [Mav21]. Clam [McM28]. Climates [Hun21]. Co [Lea25, Lea26]. Co-Operative [Lea25, Lea26]. Coarse [Sny23]. Coast [Ear26, LeC26]. Cold [Bul23]. College [Cob21]. Colloidal [Har23d]. Colorado [Ano25f]. Columbia [Dym28, Mun28]. Comes [Tho25]. Comments [Sur25]. Commercial [Mat23, Mol20, Whe23]. Commission [Tho28].

Commonwealth [But24]. Comparative [EG24]. Concerning [Emb22a, Hen20, Bea23]. Conditions [Osb21b, Osb23, Reu20]. Conducting [Orr20]. Confirmation [Bar24]. Conn. [Bin27]. Connecticut [Ano27f, Bar24, Bin27]. Conservation [Ein28, Eve23, Foe28, Fos23, Mer24, Mol20, O’M28]. Considerations [Bea23]. Constitution [Ano20b, Ano21b, Ano22b, Ano23b, Ano24c, Ano25b, Ano26b, Ano27b, Ano28c, Ano29c]. Consumer [Mat23, Mil21]. Consumption [Mil21, Pau24]. Contributions [McC29]. Control [Emb24, Reu20, Vio27]. Cooke [Nur28]. Copepods [Wil21]. County [Cle28, Tho29a]. Crustacea [Ken22]. Cultural [Cul23, Hay22, Hay24, Ken21b, Tit21]. Culture [Ano26g, Boo25, Bre23, Bul23, Can26, Dav28b, Dav29, Gor26, Lea25, Lea26, Lyd21, Lyd23, Lyd25, Moo20, Nev21, Rod24b, Sur29b, Vio28, Vio29]. Culturist [Loc29].

Dams [Ano28d]. Danger [Sur25]. date [Ano20b, Ano21b, Ano22b, Ano23b, Ano24c, Ano25b, Ano26b, Ano27b, Ano28c, Ano29c]. Debris [Til29]. Defence [Ain25]. Demonstration [Cop27, Tra21]. Denver [Ano25f]. Determination [Bar24]. Determined [Bor24]. Determining [Bel26]. Detrimental [Osb23]. Development [Cob21, Kin28, Pri23, Rod24b]. Device [Fea23b]. Diet [DL29, MBD27]. Discussion [Ano22c, Ano22f]. Disease [Bel27a, Dav26, Dav27a, Moo29]. Diseases [Ban27a, Fos25, McC29, Moo23b]. Disposal [NL28]. Distribution [Bro23b, Fea23a]. Ditches [Bak28]. Division [Rad21]. do [Pri20, Pri24]. Domestication [Kei22]. Doughty [Ain25]. Drought [Ano22c].

Early [NL28]. East [Ear26, Lai28]. Eastern [MBD27]. Economic [Ein27, Wil21]. Education [Cob28, Pri24]. EEL [Smi21b]. Effect [Ano22c, MBD27, Mer21]. Effects [NL28, Wie29]. Egg [Moo24]. Eggs [Bor27, Bro25, Har23b, Hoo29, Nor22]. Eighth [Ano28h]. Election [Ano20c, Ano25c, Ano26c, Ano27c, Ano28e, Ano29d, Ano21c, Ano22d, Ano23d]. Electricity [Cal28]. Elements [TCCM29]. Elimination [Bul27]. Enclosures [Cor20, Cor21]. Environment [Bel26, Bel29, Bro23a]. Epidemics [Bel27a]. Equipping [Orr20]. Erie [Dow23, Ken21a, Wic21]. Essential [Mol20]. Examination [Bel27b]. Experience [Nev21]. Experimental [Fie22]. Experiments [Bel27c, Bin27, Cla29, DJ24, DJ25, Dav27b, Hay21, Reu20]. Extreme [Ano22c]. Eyed [Bro25, Har23b].

Factor [MD27]. Factors [Ein27]. Facts [Ken21b]. Fairport [Dav29]. Farming [Rad26]. Fat [MBD27]. Features [Moo26b, War21b]. Federal [Fea23a]. Federation [Pol28]. Feed [ER29]. Feeding [Ano23c, Bin27, DJ25, Dav27b, Hay21]. Feeds [TCCM29]. Fertilizers [Wie29]. Fifth [Ano25f]. Fiftieth [Ano21f]. Fifty [Ano23h, Ano23i, Ano24f, Ano25f, Ano26f, Ano27f, Ano28h, Ano29a, Bow22a, Nev21, Pri21].

Fifty-Eighth [Ano28h]. Fifty-Fifth [Ano25f]. Fifty-First [Bow22a]. Fifty-Fourth [Ano24f]. Fifty-Ninth [Ano29a]. Fifty-Second [Ano23h]. Fifty-Seventh [Ano27f]. Fifty-Sixth [Ano26f]. Fifty-Third [Ano23i]. Fighting [Tra21]. Fin [Dav28a]. Fingerling [Wal25]. Fingerlings [Bin27, Dav27b, EH25, Hay26, Hes29]. First [Bow22a]. Fish [Ano22c, Ano26g, Ano28d, Ano28i, Bak28, Ban27a, Bel26, Bel27a, Bel27c, Bel28, Bel29, Bin27, Bir29, Boo25, Bre23, Bro23a, Can26, Cul21, Cul23, Dav29, Ear26, Emb22b, Fea22, Fea23a, Fea23b, Fos22, Gor26, Har23d, Hay20, Hay22, Hay24, Hen20, Jam28, Joh29, Ken21b, LeC26, LeC28, Lea25, Lea26, Loc28, Loc29, Lor27, Lyd21, Lyd23, Mad26, McC29, Mer21, Mer24, Mil21, Mol20, Moo20, Nev21, Orr20, Osb21b, Rod24b, Sny23, Sur29b, Sur29a, Syn28, Tay21, TW23, Tit21, Tit22, Vio28, Vio29, Wil20]. Fish-Cultural [Cul23, Hay22, Ken21b, Tit21]. Fish-Culture [Gor26]. Fisheries [Ano20e, Ano21f, Ano23h, Ano23i, Ano24f, Ano25f, Ano26f, Ano27f, Ano28h, Ano29a, Bal26, BMK24, Bow22a, Cal28, Cob21, Cob23b, Cob28, Lee23, Mat23, May27, Osb21a, Pop23, Pri23, Rad29, Tho28, Van29, Vio27, Whe23, Wil20, Coo29, Cul23, Rad21]. Fishermen [Pri24]. Fishery [Bab21, Eve23, McM28, Pri21, Rad21]. Fishes [Ano23c, Bro23b, Cul26, Ken22, Kid27, Leb21, Mel27, Nur28, Smi26, Tho29a, Tif21, Til29, Tit26]. Fishing [Hat25, Mun28]. Fishway [Cob27]. Flat [Moo25]. Flat-Worm [Moo25]. Flatworm [Moo26a]. Flood [Vio27]. Florida [Wil20]. Food [Bir29, Cle28, Cul26, Emb22b, Hay27, Hen20, Jam28, Joh29, Ken22, Kid27, Leb21, Met28, Met29, Mil21, Nor22, Nur28, Syn28, Til29, Tit27, TK21, Wic21]. Foods [Ano23c, DJ24, EG24, Mel27]. Forage [Dav28b, Lor27]. Forest [Mer21, Din27]. Forests [Hat25, Loc28, Tit26]. Forty [Ano20e]. Forty-Ninth [Ano20e]. Fourth [Ano24f]. Fraser [Bab21]. Fresh [Can23, DL29, Ken22, Reu20, Smi26, Tit26, Tit27]. Fresh-Water [Reu20, Smi26, Tit27]. Freshwater [Cor20, Cor21, Ell29, Smi21b]. Fry [Hes29, Nor22, Rod24a]. Function [Moo25]. Functions [Rad21]. Fundy [Mav21]. Further [Cor21, Dav27a, Moo26a, Rob22a]. Future [Cob23b, Hen23, Sur29b].

Gambusia [BA21]. Game [Mer21, Osb21b, Syn28, Tif21]. General [Ano22f, Hay26]. Georgia [Bul23]. Gill [Dav26, Dav27a, Lyd26a]. Gillworm [Lai27]. Gizzard [Tif21]. Glacier [Bro25]. Glaucous [Mun28]. Glaucous-Winged [Mun28]. Goals [McC27]. Government [Fea23a]. Grayling [Lai28]. Great [Bab21, Rad29, Van29]. Greatest [Tul20]. Green [May28]. Growth [MBD27, Rob22b, Tit22, TCCM28]. Gulf [Mav21]. Gull [Mun28]. Gyrodactylus [Emb24].

H [MD27]. Habits [Bee24, Cra21]. Halibut [Tho28]. Hartford [Ano27f]. Hatcheries [Ban27a, Tit22]. Hatchery [Bin27, Moo29, Hay20, Hay24, May28]. Hatching [Bor27, Cra21]. Headquarters [Ano20e]. Headwaters [Har23a]. Heat [Ano22c]. Her [Wil20]. Herring [Ein28]. High [Ano28d, Emb22a]. History [Ban27b, Wil21]. Hotel [Ano20e]. Hydro [Cal28]. Hydro-Electricity [Cal28].

Illinois [Tho29a]. Importance [Hen24, Moo20]. Impounded [LeC28]. Incidence [Moo29]. Incorporation [Ano20a, Ano21a, Ano22a, Ano23a, Ano24b, Ano25a, Ano26a, Ano27a, Ano28b, Ano29b]. Increase [Bro23a, Tul26]. Increased [Mil21]. Indicator [Bel29]. Industries [Rad21]. Industry [Gor26, Kin28, Pau24, Tul23]. Inhibition [Moo25]. Inland [Har23c, Hen20, Web29b]. Inorganic [TCCM29]. Inspection [Mil21]. Interests [Mun28]. International [Pol28, Smi21a, Tho28]. Intestinal [Dav24, Moo23a]. Introduction [Bro25, Hay20]. Investigation [Pol28]. Investigations [BMK24, Dav29, Sur23, TW23]. Involved [Tay21]. Irrigating [Bak28]. Irrigation [Pri23]. Isolated [Bro25].

Jay [Nur28]. Jersey [Hay20, Hay24].

Kansas [Doz25]. Kind [Pri24]. Ky [Ano20e].

Lake [Hoo29, Nur29, Coo29, Dow23, Ken21a, Wic21]. Lakes [Bel27b, Bir23, Bro26, Rod24a, Sur25, Web29b, Rad29, Van29]. Landlocked [Kei22]. Large [Bal26, Hes29, TK21]. Large-mouth [TK21]. Large-Mouthed [Hes29, Bal26]. Larvae [Joh29]. Larval [Leb21]. Laws [Ano25b, Ano26b, Ano27b, Ano28c, Ano29c]. League [Fos23]. Life [Ano22c, Ban27b, Bel28, Bro23a, Hay20, Mer21, Mer24, Sur29a]. Lines [TW23]. Liquor [NL28]. List [Ano20c, Ano21c, Ano22d, Ano23d, Ano25c, Ano26c, Ano27c, Ano28e, Ano29d, Nur28]. Live [Fea22, Fea23b]. Liver [Tho29b]. Lobster [Cra21]. Location [Tit22]. Louis [Ano23i]. Louisiana [Tul20, Tul23, Tul26, Vio26, Vio27, Vio28, Vio29]. Louisville [Ano20e].

Madison [Ano23h]. Maine [Mav21]. Maintaining [Dow23]. Man [Osb21a]. Marcos [Bal26]. Marine [Bre23]. Market [Orr20]. Maskinonge [Cha23]. Massachusetts [BMK24]. Material [Til29]. Matter [MBD27]. Means [Mil21, NL28]. Meat [DL29]. Meeting [Ano20e, Ano20d, Ano21f, Ano21e, Ano22g, Ano23h, Ano23i, Ano23g, Ano24f, Ano24e, Ano25f, Ano25e, Ano26f, Ano26e, Ano27f, Ano27e, Ano28h, Ano28g, Ano29f, Ano29a, Bow22a]. Members [Ano20c, Ano21c, Ano22d, Ano23d, Ano24a, Ano25c, Ano26c, Ano27c, Ano28e, Ano29d]. Membership [Ano20c, Ano25c, Ano26c, Ano27c, Ano28e, Ano29d, Ano21c, Ano22d, Ano23d]. Method [Bar24, Bel26, Bel27b, Fea22]. Methods [Mad26, NL28, Pop23, Sur25, Tra21]. Michigan [Lyd23, Met28, Met29, Wes29]. Migrating [Cob23c]. Migration [War21b]. Milk [Emb22b]. Mineral [MBD27]. Minneapolis [Ano29a]. Minnesota [Ano29a, Cob23a, Kid27, Sur23, Sur25, Sur29b]. Minnow [Lor27]. Minnows [Ban29, Dav28b]. Mississippi [Cul23, Sur29a, Vio27]. Missouri [Ano23i]. Mobile [Ano26f]. Modern [Mad26]. Mortality [MBD27]. Most [Vic23]. Mountain [Din27]. mouth [TK21, Wic21]. Mouthed [Bee24, Hes29, Bal26, Lyd26b]. Movements [Bel29]. Museum [Pop23]. Muskellunge [Web29a]. Mussel [Can23, Cul26]. Mussels [Cor20, Cor21, Ell29, Reu20].

National [Bro25, Din27, Hat25, Loc28, Tol25]. Natural [Emb22b, EG24, Mel27, Osb23, Rod24a, Web29b]. Need [Pri24]. Neglected [Fos23]. Nets [TW23]. Netting [Sny23]. Ninth [Ano20e, Ano29a]. Nomenclature [Cha23]. Northern [Cob23a]. Notes [Bro23b, Cor21, Cra21, Emb24, Fos25, Hay22, Lor27, Lyd23, Rad21, Tit21]. Nurseries [Rod24a]. Nutrition [MD27, McC27]. Nutritional [TCCM28, TCCM29].

Observation [Rob22b]. Observations [Bal26, Boo25, Bor25, Dav27a, Joh29, Moo26a, Smi21b]. October [Ano20e]. Octomitus [Moo23a, Moo24]. Officers [Ano21d, Ano22e, Ano23e, Ano23f, Ano24d, Ano25d, Ano26d, Ano27d, Ano28f, Ano29e]. Ohio [Ban27a, Osb21b, Tra21, TK21]. Oneida [Cle28]. Ontario [HR29, Ken21a]. Open [Ano22c, Can26]. Operations [Cul21]. Operative [Lea25, Lea26]. Other [LeC28]. Ottawa [Ano21f]. Our [Ein28, Hun21, Tol25]. Over-Stocking [Sur25]. Oxygen [BA21, Cop27]. Oyster [Kin28]. Oysters [Vic23, Wel21].

Pacific [Cob23b, Cob23c, Eve23, Kin28, McM28, Pol28, War21a]. Pack [Fie22]. Panulirus [Cra21]. Parasite [Moo23a]. Parasites [Ban29]. Parasitic [Moo25]. Parent [Rob22a]. Park [Bro25, Nur28]. Parks [Tol25]. Part [Foe28]. Past [Hen23, Sur29b]. Pathological [Hay22]. Pathology [Fos22]. Penaeus [Tul20]. Pennsylvania [Bow22a, Bul27, But24]. Perch [Cob23a, Lyd26a]. Period [Kei23]. Perpetuating [Ein28, May27]. Pike [Cob23a, Ken21a]. Pike-Perch [Cob23a]. Pimephales [Lor27]. Places [Ano20d, Ano21e, Ano22g, Ano23g, Ano24e, Ano25e, Ano26e, Ano27e, Ano28g, Ano29f]. Plan [Hay26]. Plankton [Bir23, Wie29]. Plant [Hay20, Nur28, Nur29, TCCM29, Til29]. Planting [Bro25, Har23b, Mad26]. Plants [BA21, Moo20, Tif21]. Plymouth [Leb21]. Polluted [Pow21]. Pollution [Ano22f, Bel27c, Bul27, Fie26, Har23c, Nig27, Tra21, Vio26]. Pond [Bro23b, Dav29, Moo20]. Pond-Fish [Dav29]. Ponds [BA21, Hay20, Lyd26a, Sur29a, Van25]. Populated [Hay26]. Possible [Pow21]. Post [Leb21]. Post-Larval [Leb21]. Potential [Hen24]. Poverty [Fie26]. Power [Fos23]. Practical [Fea23b, Lyd21, Osb21a, War21b]. Practices [Hay24, Ken21b]. Preliminary [Cob27, HR29, Har23d, Sco28]. Present [Hen23, Sur29b, Tul23]. Preservation [Bel20, Tay21, TW23]. Presidents’ [Ano26e, Ano27e, Ano28g, Ano29f, Ano20d, Ano21e, Ano22g, Ano23g, Ano24e, Ano25e]. Prevention [NL28]. Previously [Ken21b]. Pride [Ear26]. Prime [Mol20]. Principles [Tay21]. Priority [Cha23]. Problem [Bre23, Pau24]. Problems [Bul23, Hay22, Mat23, May27, Nig27, Van29]. Proceedings [Ano20e, Ano21f, Ano23h, Ano23i, Ano24f, Ano25f, Ano26f, Ano27f, Ano28h, Bow22a]. Producer [Mat23]. Production [Bal26, Can23, Hay27, Pau24, Tul20, Wie29]. Productivity [Bro23a]. products [TCCM29]. Program [Bul27]. Progress [Bin27, DJ25, Fie26, Lyd21, McC29, Moo23b, Moo27, Wes29]. Project [Bab28]. promelas [Lor27]. Proof [Rob22a]. Propagation [Ano28i, Bee24, Bro23b, Cob23a, Cul26, Ell29, Emb22b, Foe28, Lyd26b, May27, Web29b, Web29a, Wel21, Wes29]. Protecting [Cob23c]. Protection [Hen20, Mer21]. Protective [Syn28]. Protein [MBD27]. Proteins [TCCM29]. Proteocephalus [Ban27b, Moo26a]. Protozoa [Dav24]. Public [Bow22b, Fie26].

Quebec [Ano24f]. Question [Cha23].

Rainbow [Fos25, Ken21c, Met28]. Raising [Cor20, Cor21, Lyd26a]. Rapid [Bel27b]. Rates [TCCM28]. Razor [McM28]. Rearing [Cla29, Ein27, EH25, Reu20, Van25, Wal25]. Reclamation [Bab28, Cul26]. Reference [Bel27c, Hay22, Kid27]. Refuge [Sur29a]. Region [Coo29]. Regulation [Fie26]. Relation [BA21, Ken21b, Loc29, Mun28, Pop23, Tif21, Tit26, Vio27]. Relationship [Ken21a]. Relative [TCCM29]. Remote [Bro25]. Replenishing [Tit27]. Report [Bin27, Cob27, DJ25, Har23d, Moo23b]. Requirements [TCCM28, TCCM29]. Rescue [Cul21]. Research [Gor26]. Resources [Eve23]. Respiratory [Bel29]. Restocking [Sur25]. Restore [Bab21]. Results [Dav27b, Hay20]. Retail [Orr20]. Rival [Tho29b]. River [May28, Bab21, Bar24]. Rock [LeC26]. Rot [Dav28a]. Roughage [TCCM29]. Running [Bor27].

Salmon [Bab28, Cal28, Cob23c, Ein27, Fie22, Foe28, Har23b, Kei22, May27, NL28, Nor22, O’M28, Pol28, Pri20, War21a, War21b]. salmonis [Moo23a, Moo24]. Salmonoids [Kei23]. Salt [Ein27, Tay21, Tul26]. Salt-Water [Ein27]. Same [Lyd26a]. San [Bal26]. Sanitary [But24]. sapidissima [Bor24]. Sardines [Bea23]. Scale [Bar24]. Scales [Bor24]. Scientific [Gor26]. Scientist [Osb21a]. Scotchman [Ain25]. Scotty [Tho25]. Screens [Bak28]. Sea [Pri20]. Seafood [Vic23]. Seal [Smi21a]. Season [Lyd26a]. Seattle [Ano28h]. Second [Ano23h]. Sections [Hay26]. Seelbach [Ano20e]. Selected [EH25]. September [Ano21f, Ano23h, Ano23i, Ano24f, Ano26f, Ano29a, Bow22a]. Service [Ano20d, Ano21e, Ano22g, Ano23g, Ano24e, Ano25e, Ano26e, Ano27e, Ano28g, Ano29f]. setiferus [Tul20]. Seventh [Ano27f]. Shad [Bar24, Bor24, Ear26, Tif21]. Shall [ER29]. Shipping [Fea22]. Showing [Ano20c, Ano21c, Ano22d, Ano23d, Ano25c, Ano26c, Ano27c, Ano28e, Ano29d]. Shrimp [Tit27, Tul20, Tul23, Tul26]. Siam [Smi26]. Significance [Kei23, War21b]. Sixth [Ano26f]. Sketch [Rod24b]. Sloughs [Sur29a]. Small [BA21, Bee24, Bro26, Lyd26b, Wic21]. Small-mouth [Wic21]. Small-Mouthed [Bee24, Lyd26b]. Smallmouth [Moo25]. Smolt [Kei23]. So-called [Ken21a]. Society [Ano20e, Ano21f, Ano23h, Ano23i, Ano24f, Ano25f, Ano26f, Ano27f, Ano28h, Bow22a, Pop23, Ano29a]. Sockeye [Foe28, War21b]. Sockeyes [Rob22b]. Solution [Nig27]. Some [Bal26, Bea23, Boo25, Cla29, DJ24, DJ25, Dav27b, HR29, Hay22, Hay24, Joh29, Ken21b, Moo26b, Nur28, Orr20, Rob22b, Smi21b, Smi26, Tit21, TK21, Van29, War21b]. Sound [Leb21]. Spawning [Cra21, Moo25]. Special [Hay22, Kid27]. Species [Moo23a]. Speech [Moo28]. Spiny [Cra21]. Sportsmen [Fos23]. Spring [Bul23]. St [Ano23i]. Stage [Wal25]. Stages [NL28]. State [Bin27, Hay20, Hay24, Cle28, McM28, Moo26b, Moo27, Nig27]. States [Bab21]. Station [Din27, Bal26, Bul23, Dav29]. Status [Rad27, Rad29, Tul23]. Steelhead [Ken21c]. Stocking [Bel26, Bro23a, Bro26, Hay26, Sur25]. Stream [Moo26b, Rob22a]. Streams [Ano22f, Bel27b, Cle28, Har23c, Sur25, Tit27]. Striped [LeC26, Sco28]. Studies [Sco28]. Study [EG24, HR29, McC29, Moo23b]. Studying [Bel26]. Sturgeon [Bor25]. sturio [Bor25]. Substitutes [DL29]. Suggestions [Orr20]. Suitability [Bel26]. Sulphite [NL28]. Sulphur [Har23d]. Summary [Coo29]. Summer [Ano22c]. Superior [Coo29]. Supplies [Moo29]. Supply [BA21, Dow23, Loc28, Mer24]. Support [BA21]. Survey [McC29, Moo26b, Moo27, Osb21b, Tho29a]. Surveys [Sur23]. Symposium [Ano26g, Ano28i].

Taking [Hoo29]. Temperature [Bel28]. Temperatures [Emb22a]. Terms [Ano20d, Ano21e, Ano22g, Ano23g, Ano24e, Ano25e, Ano26e, Ano27e, Ano28g, Ano29f, TCCM29]. Test [Cop27]. Texas [Bal26]. their [Bir29, Joh29, Ken21b, Nig27, War21b]. Theory [Rob22a]. Thickly [Hay26]. Third [Ano23i]. Topminnow [BA21]. Toxic [Bel29]. Toxicity [Bel27c, Har23d]. Trade [Bel27c]. Transactions [Ano29a]. Transmission [Moo24]. Transporting [Fea23b]. Trap [Hes29]. Treating [Lai27]. Tropical [Gor26]. Troughs [Reu20]. Trout [Bin27, Bro25, Cle28, Dav24, DJ24, DJ25, Dav26, Dav27a, Dav27b, Dav28a, DL29, Din27, Doz25, Dym28, Emb22a, EG24, Emb24, EH25, Fos25, HR29, Har23b, Hat25, Hay20, Hay21, Hoo29, Ken21c, Lyd26a, MBD27, MD27, Met28, Met29, Moo23a, Moo23b, Moo24, Tit27, TCCM28, TCCM29, Wal25, Wes29].

U.S. [Rad21]. Undertaken [Moo26b]. Undeveloped [Fos23]. United [Bab21]. Unrecognized [Ken21b]. upon [Bin27, MBD27, Tho28]. Upper [Sur29a]. Use [DL29, Lai27, Lor27, Emb22b]. Used [Sur25]. Usefulness [Joh29]. Utility [Pow21]. Utilization [Sur29a].

Valley [Cul23, Vio27]. Valuable [Vic23]. Value [TCCM29]. Variations [MBD27]. Various [Wie29]. versus [Sny23]. Virginia [Lee23]. Vitamins [DJ24, MBD27, TCCM29]. vs [Bro23a].

Washington [Ano28h, May27, McM28, Nig27]. Waste [Bel27c, NL28]. Wastes [Emb22b]. Water [Bel28, Bor27, But24, Can23, Can26, Ein27, Emb22a, Ken22, Mav21, Mer24, Moo29, Rad26, Reu20, Smi26, Tit26, Tit27, Tul26, Vio26]. Waters [Bel26, Bro25, Bul27, Fie26, HR29, Hen20, Hun21, LeC28, Pow21, Sur25, Tul26, TK21]. Wells [Hen24]. West [Loc28, Pri23]. Western [Syn28]. While [Hay26]. White [Din27]. Whitefish [Dow23]. Wild [Sur29a]. Wilson [Bor24]. Winged [Mun28]. Wisconsin [Ano23h]. without [Bor27]. Word [Ain25, Wil20]. Work [Cob27, Coo29, Cul23, May27, Osb21a, Syn28, Tho28]. World [Vic23]. Worm [Moo25]. Worth [Hay26].

Year [Ano20c, Ano25c, Ano26c, Ano27c, Ano28e, Ano29d, Nev21, Ano21c, Ano22d, Ano23d]. Years [Pri21]. Yellow [Ken21a]. York [Cle28, Moo26b, Moo27]. Young [Hay27, Ken22, Rob22b, TK21, Wic21].

Zonite [Lai27].

References

Ainsworth:1925:WDD

[Ain25] S. M. Ainsworth. A word in defence of the doughty Scotchman. Transactions of the American Fisheries Society, 55(1):90–91, January 1925. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1920:CI

[Ano20a] Anonymous. Certificate of incorporation. Transactions of the American Fisheries Society, 49(4):169, September 1920. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1920:CAD

[Ano20b] Anonymous. Constitution: (as amended to date). Transactions of the American Fisheries Society, 49(4):187–190, September 1920. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1920:LMS

[Ano20c] Anonymous. List of members, 1919–1920: Showing year of election to membership. Transactions of the American Fisheries Society, 49 (4):171–185, September 1920. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1920:PTS

[Ano20d] Anonymous. Presidents, terms of service and places of meeting. Transactions of the American Fisheries Society, 49(4):170, September 1920. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1920:PAF

[Ano20e] Anonymous. Proceedings of the American Fisheries Society Forty-Ninth Annual Meeting: Louisville, Ky., October 8th, 9th and 10th, 1919 Headquarters, Seelbach Hotel. Transactions of the American Fisheries Society, 49(2):47–71, March 1920. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1921:CI

[Ano21a] Anonymous. Certificate of incorporation. Transactions of the American Fisheries Society, 50(1):4, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1921:CAD

[Ano21b] Anonymous. Constitution: (as amended to date). Transactions of the American Fisheries Society, 50(1):443–446, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1921:LMS

[Ano21c] Anonymous. List of members, 1920–1921: (showing year of election to membership). Transactions of the American Fisheries Society, 50(1):427–442, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1921:O

[Ano21d] Anonymous. Officers for 1920–1921. Transactions of the American Fisheries Society, 50(1):2, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1921:PTS

[Ano21e] Anonymous. Presidents, terms of service and places of meeting. Transactions of the American Fisheries Society, 50(1):3, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1921:PAF

[Ano21f] Anonymous. Proceedings of the American Fisheries Society: Fiftieth Annual Meeting: Ottawa, Canada: September 20, 21, 22, 1920. Transactions of the American Fisheries Society, 50(1):7–40, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1922:CI

[Ano22a] Anonymous. Certificate of incorporation. Transactions of the American Fisheries Society, 51(1):5, January 1922. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1922:CAD

[Ano22b] Anonymous. Constitution: (as amended to date). Transactions of the American Fisheries Society, 51(1):150–153, January 1922. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1922:EDE

[Ano22c] Anonymous. Effect of drought and extreme heat of summer on fish life: Open discussion. Transactions of the American Fisheries Society, 51(1):133–135, January 1922. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1922:LMS

[Ano22d] Anonymous. List of members, 1921–1922: (showing year of election to membership). Transactions of the American Fisheries Society, 51(1):136–149, January 1922. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1922:O

[Ano22e] Anonymous. Officers for 1921–1922. Transactions of the American Fisheries Society, 51(1):3, January 1922. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1922:PSG

[Ano22f] Anonymous. Pollution of streams: A general discussion. Transactions of the American Fisheries Society, 51(1):126–132, January 1922. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1922:PTS

[Ano22g] Anonymous. Presidents, terms of service and places of meeting. Transactions of the American Fisheries Society, 51(1):4, January 1922. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1923:CI

[Ano23a] Anonymous. Certificate of incorporation. Transactions of the American Fisheries Society, 52(1):5, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1923:CAD

[Ano23b] Anonymous. Constitution: (as amended to date). Transactions of the American Fisheries Society, 52(1):254–257, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1923:FFF

[Ano23c] Anonymous. Foods and feeding of fishes. Transactions of the American Fisheries Society, 53(1):137–159, January 1923. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1923:LMS

[Ano23d] Anonymous. List of members, 1922–1923: (showing year of election to membership.). Transactions of the American Fisheries Society, 52(1):239–253, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1923:Oa

[Ano23e] Anonymous. Officers for 1922–1923. Transactions of the American Fisheries Society, 52(1):3, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1923:Ob

[Ano23f] Anonymous. Officers for 1923–1924. Transactions of the American Fisheries Society, 53(1):3, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1923:PTS

[Ano23g] Anonymous. Presidents, terms of service and places of meeting. Transactions of the American Fisheries Society, 52(1):4, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1923:PAFa

[Ano23h] Anonymous. Proceedings of the American Fisheries Society: Fifty-Second Annual Meeting at Madison, Wisconsin: September 6, 7, 8, 1922. Transactions of the American Fisheries Society, 52(1):7–30, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1923:PAFb

[Ano23i] Anonymous. Proceedings of the American Fisheries Society: Fifty-Third Annual Meeting at St. Louis, Missouri: September 17, 18, 19, 1923. Transactions of the American Fisheries Society, 53(1): 5–22, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1924:AM

[Ano24a] Anonymous. Active members. Transactions of the American Fisheries Society, 54(1):207–218, January 1924. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1924:CI

[Ano24b] Anonymous. Certificate of incorporation. Transactions of the American Fisheries Society, 54(1):203, January 1924. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1924:CAD

[Ano24c] Anonymous. Constitution: (as amended to date). Transactions of the American Fisheries Society, 54(1):204–206, January 1924. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1924:O

[Ano24d] Anonymous. Officers for 1924–1925. Transactions of the American Fisheries Society, 54(1):3, January 1924. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1924:PTS

[Ano24e] Anonymous. Presidents, terms of service and places of meeting. Transactions of the American Fisheries Society, 54(1):4, January 1924. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1924:PAF

[Ano24f] Anonymous. Proceedings of the American Fisheries Society: Fifty-Fourth Annual Meeting at Quebec, Canada: September 10, 11, 12, 1924. Transactions of the American Fisheries Society, 54(1):7–21, January 1924. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1925:CI

[Ano25a] Anonymous. Certificate of incorporation. Transactions of the American Fisheries Society, 55(1):193, January 1925. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1925:CLA

[Ano25b] Anonymous. Constitution and by-laws. (as amended to date). Transactions of the American Fisheries Society, 55(1):194–197, January 1925. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1925:LMS

[Ano25c] Anonymous. List of members, 1925–1926. (showing year of election to membership). Transactions of the American Fisheries Society, 55(1):198–212, January 1925. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1925:O

[Ano25d] Anonymous. Officers for 1925–1926. Transactions of the American Fisheries Society, 55(1):3, January 1925. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1925:PTS

[Ano25e] Anonymous. Presidents terms of service and places of meeting. Transactions of the American Fisheries Society, 55(1):4, January 1925. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1925:PAF

[Ano25f] Anonymous. Proceedings of the American Fisheries Society: Fifty-Fifth Annual Meeting at Denver Colorado: August 17, 18, 19, 1925. Transactions of the American Fisheries Society, 55(1):7–29, January 1925. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1926:CI

[Ano26a] Anonymous. Certificate of incorporation. Transactions of the American Fisheries Society, 56(1):241, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1926:CLA

[Ano26b] Anonymous. Constitution and by-laws. (as amended to date). Transactions of the American Fisheries Society, 56(1):242–245, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1926:LMS

[Ano26c] Anonymous. List of members, 1926–1927. (showing year of election to membership). Transactions of the American Fisheries Society, 56(1):246–262, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1926:O

[Ano26d] Anonymous. Officers for 1926–1927. Transactions of the American Fisheries Society, 56(1):3, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1926:PTS

[Ano26e] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 56(1):4–5, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1926:PAF

[Ano26f] Anonymous. Proceedings of the American Fisheries Society: Fifty-Sixth Annual Meeting at Mobile, Alabama: September 22, 23, 24, 1926. Transactions of the American Fisheries Society, 56(1):9–40, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1926:SFC

[Ano26g] Anonymous. Symposium on fish culture. Transactions of the American Fisheries Society, 56(1):130–155, January 1926. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1927:CI

[Ano27a] Anonymous. Certificate of incorporation. Transactions of the American Fisheries Society, 57(1):303, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1927:CLA

[Ano27b] Anonymous. Constitution and by-laws. (as amended to date). Transactions of the American Fisheries Society, 57(1):304–307, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1927:LMS

[Ano27c] Anonymous. List of members, 1927–1928. (showing year of election to membership). Transactions of the American Fisheries Society, 57(1):308–324, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1927:O

[Ano27d] Anonymous. Officers for 1927–1928. Transactions of the American Fisheries Society, 57(1):3, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1927:PTS

[Ano27e] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 57(1):4–5, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1927:PAF

[Ano27f] Anonymous. Proceedings of the American Fisheries Society: Fifty-Seventh Annual Meeting at Hartford, Connecticut: August 8, 9, 10, 1927. Transactions of the American Fisheries Society, 57(1):9–36, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1928:B

[Ano28a] Anonymous. Banquet. Transactions of the American Fisheries Society, 58(1):245, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1928:CI

[Ano28b] Anonymous. Certificate of incorporation. Transactions of the American Fisheries Society, 58(1):249–251, January 1928. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1928:CLA

[Ano28c] Anonymous. Constitution and by-laws. (as amended to date). Transactions of the American Fisheries Society, 58(1):252–255, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1928:HDF

[Ano28d] Anonymous. High dams and fish. Transactions of the American Fisheries Society, 58(1):114–116, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1928:LMS

[Ano28e] Anonymous. List of members, 1928–1929: (showing year of election to membership). Transactions of the American Fisheries Society, 58(1):256–274, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1928:O

[Ano28f] Anonymous. Officers for 1928–1929. Transactions of the American Fisheries Society, 58(1):3, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1928:PTS

[Ano28g] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 58(1):4–5, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1928:PAF

[Ano28h] Anonymous. Proceedings of the American Fisheries Society: Fifty-Eighth Annual Meeting at Seattle, Washington: August 29, 30 and 31, 1928. Transactions of the American Fisheries Society, 58(1): 9–35, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1928:SFP

[Ano28i] Anonymous. Symposium on fish propagation. Transactions of the American Fisheries Society, 58(1):198–200, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1929:TAF

[Ano29a] Anonymous. Transactions of the American Fisheries Society: Fifty-Ninth Annual Meeting at Minneapolis, Minnesota: September 9, 10, and 11, 1929. Transactions of the American Fisheries Society, 59 (1):9–41, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1929:CI

[Ano29b] Anonymous. Certificate of incorporation. Transactions of the American Fisheries Society, 59(1):275, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1929:CLA

[Ano29c] Anonymous. Constitution and by-laws: (as amended to date). Transactions of the American Fisheries Society, 59(1):276–279, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1929:LMS

[Ano29d] Anonymous. List of members, 1929–1930: (showing year of election to membership). Transactions of the American Fisheries Society, 59(1):280–297, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1929:O

[Ano29e] Anonymous. Officers for 1929–1930. Transactions of the American Fisheries Society, 59(1):3, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1929:PTS

[Ano29f] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 59(1):4–5, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barney:1921:RCA

[BA21] R. L. Barney and B. J. Anson. Relation of certain aquatic plants to oxygen supply and to capacity of small ponds to support the topminnow (Gambusia affinis). Transactions of the American Fisheries Society, 50(1):268–278, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Babcock:1921:CUS

[Bab21] John Pease Babcock. Canada and the United States can restore the Great Fraser River fishery. Transactions of the American Fisheries Society, 50(1):232–246, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Babcock:1928:CAS

[Bab28] John Pease Babcock. A Canadian-American salmon reclamation project. Transactions of the American Fisheries Society, 58(1):44– 51, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baker:1928:FSI

[Bak28] Shirley Baker. Fish screens in irrigating ditches. Transactions of the American Fisheries Society, 58(1):80–82, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baldwin:1926:SOP

[Bal26] O. N. Baldwin. Some observations on the production of largemouthed black bass at San Marcos, Texas, Fisheries Station. Transactions of the American Fisheries Society, 56(1):50–52, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bangham:1927:DFO

[Ban27a] Ralph H. Bangham. Diseases of fish in Ohio hatcheries. Transactions of the American Fisheries Society, 57(1):223–230, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bangham:1927:LHB

[Ban27b] Ralph H. Bangham. Life history of bass cestode Proteocephalus ambloplitis. Transactions of the American Fisheries Society, 57(1): 206–209, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bangham:1929:PBM

[Ban29] R. V. Bangham. Parasites of bait minnows. Transactions of the American Fisheries Society, 59(1):198–201, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barney:1924:CBS

[Bar24] R. L. Barney. A confirmation of Borodin’s scale method of age determination of Connecticut River shad. Transactions of the American Fisheries Society, 54(1):168–177, January 1924. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beard:1923:SCC

[Bea23] Harry R. Beard. Some considerations concerning the canning of sardines. Transactions of the American Fisheries Society, 53(1): 122–130, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beeman:1924:HPS

[Bee24] Henry W. Beeman. Habits and propagation of the small-mouthed black bass. Transactions of the American Fisheries Society, 54(1): 92–107, January 1924. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1920:PA

[Bel20] David L. Belding. The preservation of the alewife. Transactions of the American Fisheries Society, 49(2):92–104, March 1920. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1926:NMS

[Bel26] David L. Belding. A new method of studying fish environment and determining the suitability of waters for stocking. Transactions of the American Fisheries Society, 56(1):79–82, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1927:FDE

[Bel27a] David L. Belding. Fish disease epidemics. Transactions of the American Fisheries Society, 57(1):217–222, January 1927. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1927:RME

[Bel27b] David L. Belding. A rapid method for the examination of lakes and streams. Transactions of the American Fisheries Society, 57(1):

73–84, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1927:TEF

[Bel27c] David L. Belding. Toxicity experiments with fish in reference to trade waste pollution. Transactions of the American Fisheries Society, 57(1):100–119, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1928:WTF

[Bel28] David L. Belding. Water temperature and fish life. Transactions of the American Fisheries Society, 58(1):98–105, January 1928. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1929:RMF

[Bel29] David L. Belding. The respiratory movements of fish as an indicator of a toxic environment. Transactions of the American Fisheries Society, 59(1):238–245, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bing:1927:PRU

[Bin27] Franklin C. Bing. A progress report upon feeding experiments with brook trout fingerlings at the Connecticut State Fish Hatchery, Burlington, Conn. Transactions of the American Fisheries Society, 57(1):266–280, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Birge:1923:PL

[Bir23] E. A. Birge. The plankton of the lakes. Transactions of the American Fisheries Society, 52(1):118–130, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Birge:1929:FTF

[Bir29] E. A. Birge. Fish and their food. Transactions of the American Fisheries Society, 59(1):188–194, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belding:1924:FIM

[BMK24] David L. Belding, Arthur Merrill, and James Kitson. Fisheries investigations in Massachusetts. Transactions of the American Fisheries Society, 54(1):29–47, January 1924. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Booth:1925:SOF

[Boo25] D. C. Booth. Some observations on fish culture. Transactions of the American Fisheries Society, 55(1):161–166, January 1925. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Borodin:1924:ASA

[Bor24] N. Borodin. Age of shad (Alosa sapidissima Wilson) as determined by the scales. Transactions of the American Fisheries Society, 54(1): 178–184, January 1924. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Borodin:1925:BOA

[Bor25] N. Borodin. Biological observations on the Atlantic sturgeon (Acipenser sturio). Transactions of the American Fisheries Society, 55(1):184–190, January 1925. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Borodin:1927:HER

[Bor27] N. Borodin. Hatching eggs without running water. Transactions of the American Fisheries Society, 57(1):162–164, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bower:1922:PAF

[Bow22a] Ward T. Bower. Proceedings of the American Fisheries Society: Fifty-First Annual Meeting at Allentown, Pennsylvania, September 5, 6, 7, 1921. Transactions of the American Fisheries Society, 51 (1):7–32, January 1922. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bower:1922:PA

[Bow22b] Ward T. Bower. Public aquariums. Transactions of the American Fisheries Society, 51(1):117–125, January 1922. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Breder:1923:PMF

[Bre23] C. M. Breder, Jr. The problem of marine fish culture. Transactions of the American Fisheries Society, 52(1):210–218, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1923:AEV

[Bro23a] Ernest Clive Brown. Adjustment of environment vs. stocking — to increase the productivity of fish life. Transactions of the American Fisheries Society, 52(1):131–145, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1923:NPD

[Bro23b] G. W. N. Brown. Notes on the propagation and distribution of pond fishes. Transactions of the American Fisheries Society, 53(1): 184–190, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brooks:1925:ITP

[Bro25] James P. Brooks. Introduction of trout and the planting of eyed eggs in remote and isolated waters of Glacier National Park. Transactions of the American Fisheries Society, 55(1):149–153, January 1925. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1926:BBS

[Bro26] Dell Brown. Bream or bluegills for stocking small lakes. Transactions of the American Fisheries Society, 56(1):207–209, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bullock:1923:PBC

[Bul23] Charles A. Bullock. Problems in bass culture at the Cold Spring, Georgia Station. Transactions of the American Fisheries Society, 53(1):131–137, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buller:1927:PPE

[Bul27] Nathan R. Buller. The Pennsylvania program for the elimination of pollution of waters. Transactions of the American Fisheries Society, 57(1):39–48, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Butler:1924:SWB

[But24] Nathan R. Butler. The Sanitary Water Board of the Commonwealth of Pennsylvania. Transactions of the American Fisheries Society, 54(1):163–167, January 1924. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Calderwood:1928:HES

[Cal28] W. J. Calderwood. Hydro-electricity and salmon fisheries. Transactions of the American Fisheries Society, 58(1):154–160, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Canfield:1923:PFW

[Can23] H. L. Canfield. Production of the fresh water mussel. Transactions of the American Fisheries Society, 53(1):171–175, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Canfield:1926:OWF

[Can26] H. L. Canfield. Open water fish culture. Transactions of the American Fisheries Society, 56(1):190–194, January 1926. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chambers:1923:MQP

[Cha23] E. T. D. Chambers. The maskinonge: A question of priority in nomenclature. Transactions of the American Fisheries Society, 52 (1):171–177, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clapp:1929:SER

[Cla29] Alva Clapp. Some experiments in rearing channel catfish. Transactions of the American Fisheries Society, 59(1):114–117, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clemens:1928:FTS

[Cle28] Wilbert A. Clemens. The food of trout from the streams of Oneida County, New York State. Transactions of the American Fisheries Society, 58(1):183–197, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cobb:1921:DCF

[Cob21] John N. Cobb. Development of the College of fisheries. Transactions of the American Fisheries Society, 50(1):56–68, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cobb:1923:PPP

[Cob23a] Eben W. Cobb. Pike-perch propagation in Northern Minnesota. Transactions of the American Fisheries Society, 53(1):95–105, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cobb:1923:FFP

[Cob23b] John N. Cobb. The future of the fisheries of the Pacific. Transactions of the American Fisheries Society, 53(1):65–73, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cobb:1923:PMP

[Cob23c] John N. Cobb. Protecting migrating Pacific salmon. Transactions of the American Fisheries Society, 52(1):146–156, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cobb:1927:PRF

[Cob27] John N. Cobb. Preliminary report of fishway work. Transactions of the American Fisheries Society, 57(1):181–201, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cobb:1928:FE

[Cob28] John N. Cobb. Fisheries education. Transactions of the American Fisheries Society, 58(1):146–153, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cook:1929:BSW

[Coo29] W. A. Cook. A brief summary of work of the Bureau of Fisheries in the Lake Superior region. Transactions of the American Fisheries Society, 59(1):56–62, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Copeland:1927:DOT

[Cop27] William R. Copeland. Demonstration of the oxygen test. Transactions of the American Fisheries Society, 57(1):231–239, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Corwin:1920:RFM

[Cor20] Roy S. Corwin. Raising freshwater mussels in enclosures. Transactions of the American Fisheries Society, 49(2):81–84, March 1920. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Corwin:1921:FNR

[Cor21] Roy S. Corwin. Further notes on raising freshwater mussels in enclosures. Transactions of the American Fisheries Society, 50(1): 307–311, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Crawford:1921:SHS

[Cra21] D. R. Crawford. Spawning habits of the spiny lobster (Panulirus argus), with notes on artificial hatching. Transactions of the American Fisheries Society, 50(1):312–319, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Culler:1921:FRO

[Cul21] C. F. Culler. Fish rescue operations. Transactions of the American Fisheries Society, 50(1):247–250, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Culler:1923:FCW

[Cul23] C. F. Culler. Fish-cultural work of the Bureau of Fisheries in the Mississippi Valley. Transactions of the American Fisheries Society, 52(1):186–192, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Culler:1926:RFF

[Cul26] C. F. Culler. Reclamation of food fishes and mussel propagation. Transactions of the American Fisheries Society, 56(1):175–180, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1924:IPT

[Dav24] H. S. Davis. The intestinal protozoa of trout. Transactions of the American Fisheries Society, 54(1):57–67, January 1924. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1926:NGD

[Dav26] H. S. Davis. A new gill disease of trout. Transactions of the American Fisheries Society, 56(1):156–160, January 1926. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1927:FOG

[Dav27a] H. S. Davis. Further observations on the gill disease of trout. Transactions of the American Fisheries Society, 57(1):210–216, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1927:SRF

[Dav27b] H. S. Davis. Some results of feeding experiments with trout fingerlings. Transactions of the American Fisheries Society, 57(1):

281–287, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1928:FRT

[Dav28a] H. S. Davis. Fin rot in trout. Transactions of the American Fisheries Society, 58(1):89–91, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1928:FMB

[Dav28b] H. S. Davis. Forage minnows and bass culture. Transactions of the American Fisheries Society, 58(1):92–97, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1929:IPF

[Dav29] H. S. Davis. Investigations in pond-fish culture at the Fairport Biological Station. Transactions of the American Fisheries Society, 59 (1):86–93, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Denmead:1920:CB

[Den20] Talbot Denmead. The Chesapeake Bay. Transactions of the American Fisheries Society, 49(2):78–80, March 1920. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dinsmore:1927:BTS

[Din27] A. H. Dinsmore. A brook trout station in the White Mountain National Forest. Transactions of the American Fisheries Society, 57(1):202–205, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1924:SEA

[DJ24] H. S. Davis and M. C. James. Some experiments on the addition of vitamins to trout foods. Transactions of the American Fisheries Society, 54(1):77–91, January 1924. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1925:SET

[DJ25] H. S. Davis and M. C. James. Some experiments in trout feeding — a progress report. Transactions of the American Fisheries Society, 55(1):115–128, January 1925. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1929:USF

[DL29] H. S. Davis and R. F. Lord. The use of substitutes for fresh meat in the diet of trout. Transactions of the American Fisheries Society, 59(1):160–167, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Downing:1923:WMS

[Dow23] S. W. Downing. Are we maintaining the supply of whitefish in Lake Erie? Transactions of the American Fisheries Society, 53(1):62–64, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Doze:1925:BTK

[Doz25] J. B. Doze. The barbed trout of Kansas. Transactions of the American Fisheries Society, 55(1):167–183, January 1925. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dymond:1928:TBC

[Dym28] J. R. Dymond. The trout of British Columbia. Transactions of the American Fisheries Society, 58(1):71–77, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Earle:1926:SPF

[Ear26] Swepson Earle. The shad — pride fish of the East Coast. Transactions of the American Fisheries Society, 56(1):210–212, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Embody:1924:CSN

[EG24] G. C. Embody and Myron Gordon. A comparative study of natural and artificial foods of brook trout. Transactions of the American Fisheries Society, 54(1):185–200, January 1924. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Embody:1925:ARB

[EH25] George C. Embody and Charles O. Hayford. The advantage of rearing brook trout fingerlings from selected breeders. Transactions of the American Fisheries Society, 55(1):135–148, January 1925. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Einarsen:1927:EFS

[Ein27] Arthur S. Einarsen. Economic factors in the salt-water rearing of salmon. Transactions of the American Fisheries Society, 57(1):288– 293, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Einarsen:1928:POH

[Ein28] Arthur S. Einarsen. Perpetuating our herring — by conservation. Transactions of the American Fisheries Society, 58(1):132–138, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ellis:1929:APF

[Ell29] M. M. Ellis. The artificial propagation of freshwater mussels. Transactions of the American Fisheries Society, 59(1):217–223, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Embody:1922:CHW

[Emb22a] G. C. Embody. Concerning high water temperatures and trout. Transactions of the American Fisheries Society, 51(1):58–64, January 1922. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Embody:1922:UCM

[Emb22b] G. C. Embody. The use of certain milk wastes in the propagation of natural fish food. Transactions of the American Fisheries Society, 51 (1):76–81, January 1922. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Embody:1924:NCG

[Emb24] Geo C. Embody. Notes on the control of Gyrodactylus on trout. Transactions of the American Fisheries Society, 54(1):48–53, January 1924. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Einarsen:1929:WSW

[ER29] Arthur S. Einarsen and Loyd Royal. What shall we feed? Transactions of the American Fisheries Society, 59(1):268–271, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Evermann:1923:CFR

[Eve23] Barton Warren Evermann. The conservation of the fishery resources of the Pacific. Transactions of the American Fisheries Society, 53(1): 210–214, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fearnow:1922:NMS

[Fea22] Edgar C. Fearnow. A new method of shipping live fish. Transactions of the American Fisheries Society, 51(1):37–47, January 1922. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fearnow:1923:FDF

[Fea23a] E. C. Fearnow. Fish distribution by the federal government. Transactions of the American Fisheries Society, 53(1):160–170, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fearnow:1923:NPD

[Fea23b] Edgar C. Fearnow. A new and practical device for transporting live fish. Transactions of the American Fisheries Society, 52(1):98–117, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fiedler:1922:BAE

[Fie22] Reginald H. Fiedler. Bacteriological analysis of an experimental pack of canned salmon. Transactions of the American Fisheries Society, 51(1):110–116, January 1922. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Field:1926:PPT

[Fie26] George W. Field. Progress or poverty through regulation of pollution of public waters. Transactions of the American Fisheries Society, 56(1):97–100, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Foerster:1928:PPC

[Foe28] R. E. Foerster. Propagation’s part in the conservation of sockeye salmon. Transactions of the American Fisheries Society, 58(1):52– 67, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Foster:1922:FP

[Fos22] W. T. Foster. Fish pathology. Transactions of the American Fisheries Society, 51(1):107–109, January 1922. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Foster:1923:SLN

[Fos23] Fred J. Foster. The Sportsmen’s League — a neglected and undeveloped power for conservation. Transactions of the American Fisheries Society, 53(1):50–54, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Foster:1925:NCD

[Fos25] Fred J. Foster. Notes on certain diseases of rainbow trout. Transactions of the American Fisheries Society, 55(1):80–89, January 1925. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gordon:1926:TAF

[Gor26] Myron Gordon. Tropical aquarium fish-culture in industry and scientific research. Transactions of the American Fisheries Society, 56(1):227–237, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hare:1923:H

[Har23a] F. E. Hare. Headwaters. Transactions of the American Fisheries Society, 53(1):106–109, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harrison:1923:PES

[Har23b] C. W. Harrison. Planting eyed salmon and trout eggs. Transactions of the American Fisheries Society, 53(1):191–200, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hart:1923:PIS

[Har23c] M. D. Hart. Pollution of inland streams. Transactions of the American Fisheries Society, 52(1):166–170, January 1923. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harukawa:1923:PRT

[Har23d] Chukichi Harukawa. Preliminary report on the toxicity of colloidal sulphur to fish. Transactions of the American Fisheries Society, 52(1):219–224, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hatton:1925:TFN

[Hat25] John H. Hatton. Trout fishing in the national forests. Transactions of the American Fisheries Society, 55(1):154–160, January 1925. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayford:1920:BRI

[Hay20] Chas O. Hayford. Beneficial results of the introduction of plant life in trout ponds at the New Jersey State Fish Hatchery. Transactions of the American Fisheries Society, 49(2):85–91, March 1920. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayford:1921:TFE

[Hay21] Chas O. Hayford. Trout feeding experiments. Transactions of the American Fisheries Society, 50(1):251–256, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayford:1922:SFC

[Hay22] Charles O. Hayford. Some fish-cultural notes, with special reference to pathological problems. Transactions of the American Fisheries Society, 51(1):97–106, January 1922. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayford:1924:SFC

[Hay24] Charles O. Hayford. Some fish cultural practices in the New Jersey State Hatchery. Transactions of the American Fisheries Society, 54(1):108–137, January 1924. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayford:1926:GPS

[Hay26] Charles O. Hayford. Is the general plan of stocking with fingerlings in thickly populated sections worth while? Transactions of the American Fisheries Society, 56(1):181–189, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayford:1927:APF

[Hay27] Charles O. Hayford. Artificial production of food for young bass. Transactions of the American Fisheries Society, 57(1):143–149, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henshall:1920:CPF

[Hen20] James Alexander Henshall. Concerning the protection of fish, fish food and inland waters. Transactions of the American Fisheries Society, 49(3):141–147, June 1920. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henshall:1923:APP

[Hen23] James Alexander Henshall. Angling: Past, present, future. Transactions of the American Fisheries Society, 53(1):176–183, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henshall:1924:PIA

[Hen24] James Alexander Henshall. The potential importance of artesian wells. Transactions of the American Fisheries Society, 54(1):68–71, January 1924. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hesen:1929:TCL

[Hes29] Herman O. Hesen. A trap for the capture of large-mouthed bass fry and fingerlings. Transactions of the American Fisheries Society, 59(1):118–125, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoofnagle:1929:TCL

[Hoo29] G. W. Hoofnagle. The taking and care of lake trout eggs. Transactions of the American Fisheries Society, 59(1):53–55, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harkness:1929:PSS

[HR29] William J. K. Harkness and William E. Ricker. A preliminary study of some trout waters of Ontario. Transactions of the American Fisheries Society, 59(1):256–267, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huntsman:1921:COA

[Hun21] A. G. Huntsman. Climates of our Atlantic waters. Transactions of the American Fisheries Society, 50(1):326–333, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

James:1928:MAF

[Jam28] M. C. James. More about fish food. Transactions of the American Fisheries Society, 58(1):170–174, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1929:SOC

[Joh29] M. S. Johnson. Some observations on chironomid larvae and their usefulness as fish food. Transactions of the American Fisheries Society, 59(1):153–159, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Keil:1922:DLS

[Kei22] W. M. Keil. The domestication of landlocked salmon breeders. Transactions of the American Fisheries Society, 51(1):48–57, January 1922. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Keil:1923:BSS

[Kei23] William M. Keil. The biological significance of the smolt period in certain salmonoids. Transactions of the American Fisheries Society, 52(1):178–183, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kendall:1921:RCB

[Ken21a] William Converse Kendall. The relationship of the so-called blue pike and yellow pike of Lake Erie and Lake Ontario. Transactions of the American Fisheries Society, 50(1):257–267, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kendall:1921:SPU

[Ken21b] William Converse Kendall. Some previously unrecognized anatomical facts and their relation to fish-cultural practices. Transactions of the American Fisheries Society, 50(1):43–55, January 1921. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kendall:1921:WRT

[Ken21c] William Converse Kendall. What are rainbow trout and steelhead trout? Transactions of the American Fisheries Society, 50(1):187– 199, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kendall:1922:FWC

[Ken22] William Converse Kendall. Fresh water crustacea as food for young fishes. Transactions of the American Fisheries Society, 51(1):70–75, January 1922. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kidd:1927:FMF

[Kid27] Patience Ellis Kidd. The food of Minnesota fishes with special reference to the algae. Transactions of the American Fisheries Society, 57(1):85–91, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kincaid:1928:DOI

[Kin28] Trevor B. Kincaid. Development of oyster industry of the Pacific. Transactions of the American Fisheries Society, 58(1):117–122, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Laird:1927:UZT

[Lai27] James A. Laird. The use of zonite in treating gillworm. Transactions of the American Fisheries Society, 57(1):177–180, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Laird:1928:GE

[Lai28] James A. Laird. Grayling in the East. Transactions of the American Fisheries Society, 58(1):167–169, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leach:1925:COF

[Lea25] Glen C. Leach. Co-operative fish culture. Transactions of the American Fisheries Society, 55(1):102–114, January 1925. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leach:1926:COF

[Lea26] Glen C. Leach. Co-operative fish culture. Transactions of the American Fisheries Society, 56(1):161–168, January 1926. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lebour:1921:FLP

[Leb21] Marie V. Lebour. The food of the larval and post-larval fishes of Plymouth Sound. Transactions of the American Fisheries Society, 50(1):345–352, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

LeCompte:1926:SBR

[LeC26] E. Lee LeCompte. The striped bass (rock fish) of the Atlantic Coast. Transactions of the American Fisheries Society, 56(1):203–206, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

LeCompte:1928:BOF

[LeC28] E. Lee LeCompte. Bass and other fish in impounded waters. Transactions of the American Fisheries Society, 58(1):83–86, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lee:1923:FV

[Lee23] McDonald Lee. Fisheries in Virginia. Transactions of the American Fisheries Society, 53(1):207–209, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Locke:1928:NFW

[Loc28] S. B. Locke. The national forests of the West and the fish supply. Transactions of the American Fisheries Society, 58(1):201–204, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lockwood:1929:RBA

[Loc29] Kenneth F. Lockwood. The relation between angler and fish culturist. Transactions of the American Fisheries Society, 59(1):246–249, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lord:1927:NUB

[Lor27] Russell F. Lord, Jr. Notes on the use of the blackhead minnow, Pimephales promelas, as a forage fish. Transactions of the American Fisheries Society, 57(1):92–99, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lydell:1921:PPF

[Lyd21] Dwight Lydell. Progress in practical fish culture. Transactions of the American Fisheries Society, 50(1):221–231, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lydell:1923:BNF

[Lyd23] Dwight Lydell. Brief notes on fish culture in Michigan. Transactions of the American Fisheries Society, 52(1):184–185, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lydell:1925:BBC

[Lyd25] Dwight H. Lydell. Black bass culture. Transactions of the American Fisheries Society, 55(1):37–62, January 1925. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lydell:1926:RTB

[Lyd26a] Dwight Lydell. Raising trout in bass ponds and perch in blue gill ponds in the same season. Transactions of the American Fisheries Society, 56(1):47–49, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lydell:1926:SMB

[Lyd26b] Dwight Lydell. Small-mouthed black bass propagation. Transactions of the American Fisheries Society, 56(1):43–46, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Madsen:1926:MMF

[Mad26] David H. Madsen. Modern methods of fish planting. Transactions of the American Fisheries Society, 56(1):195–202, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matthews:1923:PCF

[Mat23] J. H. Matthews. Problems of the commercial fisheries from producer to consumer. Transactions of the American Fisheries Society, 52(1): 193–200, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mavor:1921:CWB

[Mav21] James W. Mavor. Circulation of water in Bay of Fundy and Gulf of Maine. Transactions of the American Fisheries Society, 50(1): 334–344, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mayhall:1927:PWP

[May27] L. E. Mayhall. Propagation work and problems in perpetuating the salmon fisheries of Washington. Transactions of the American Fisheries Society, 57(1):165–170, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mayhall:1928:GRH

[May28] John Mayhall. The Green River Hatchery. Transactions of the American Fisheries Society, 58(1):87–88, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCay:1927:EVV

[MBD27] C. M. McCay, F. C. Bing, and W. S. Dilley. The effect of variations in vitamins, protein, fat and mineral matter in the diet upon the growth and mortality of Eastern brook trout. Transactions of the American Fisheries Society, 57(1):240–249, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCay:1927:GN

[McC27] C. M. McCay. Goals in nutrition. Transactions of the American Fisheries Society, 57(1):261–265, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCay:1929:CBS

[McC29] C. M. McCay. Contributions of the biological survey to progress in the study of fish diseases. Transactions of the American Fisheries Society, 59(1):197, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McMillin:1928:PRC

[McM28] H. C. McMillin. The Pacific razor clam fishery in the State of Washington. Transactions of the American Fisheries Society, 58(1):106– 113, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCay:1927:FHN

[MD27] C. M. McCay and W. E. Dilley. Factor H in the nutrition of trout. Transactions of the American Fisheries Society, 57(1):250–260, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mellen:1927:NAF

[Mel27] Ida Mellen. The natural and artificial foods of fishes. Transactions of the American Fisheries Society, 57(1):120–142, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mercier:1921:FPE

[Mer21] Honore Mercier. Forest protection and its effect on fish and game life. Transactions of the American Fisheries Society, 50(1):96–105, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mercier:1924:CWS

[Mer24] Honore Mercier. Conservation of water supply as affecting fish life. Transactions of the American Fisheries Society, 54(1):72–76, January 1924. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Metzelaar:1928:FRT

[Met28] Jan Metzelaar. The food of the rainbow trout in Michigan. Transactions of the American Fisheries Society, 58(1):178–182, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Metzelaar:1929:FTM

[Met29] Jan Metzelaar. The food of the trout in Michigan. Transactions of the American Fisheries Society, 59(1):146–152, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Millett:1921:AFI

[Mil21] Arthur L. Millett. Adequate fish inspection: A means to better fish for the consumer and to increased fish food consumption. Transactions of the American Fisheries Society, 50(1):153–162, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mollan:1920:PEC

[Mol20] William K. Mollan. The prime essential in commercial fish conservation. Transactions of the American Fisheries Society, 49(2):72–77, March 1920. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moore:1920:PIP

[Moo20] Emmeline Moore. Plants of importance in pond fish culture. Transactions of the American Fisheries Society, 49(3):148–152, June 1920. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moore:1923:OSN

[Moo23a] Emmeline Moore. Octomitus salmonis, a new species of intestinal parasite in trout. Transactions of the American Fisheries Society, 52 (1):74–97, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moore:1923:RPS

[Moo23b] Emmeline Moore. A report of progress on the study of trout diseases. Transactions of the American Fisheries Society, 53(1):74–94, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moore:1924:TOS

[Moo24] Emmeline Moore. The transmission of Octomitus salmonis in the egg of trout. Transactions of the American Fisheries Society, 54(1): 54–56, January 1924. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moore:1925:ISF

[Moo25] Emmeline Moore. The inhibition of the spawning function of the smallmouth bass by a parasitic flat-worm. Transactions of the American Fisheries Society, 55(1):33–36, January 1925. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moore:1926:FOB

[Moo26a] Emmeline Moore. Further observations on the bass flatworm — Proteocephalus ambloplitis. Transactions of the American Fisheries Society, 56(1):91–96, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moore:1926:SFS

[Moo26b] Emmeline Moore. Some features of the stream survey undertaken in New York State. Transactions of the American Fisheries Society, 56(1):108–121, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moore:1927:PBS

[Moo27] Emmeline Moore. Progress of the biological survey in New York State. Transactions of the American Fisheries Society, 57(1):65–72, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moore:1928:BS

[Moo28] Emmeline Moore. Banquet speech. Transactions of the American Fisheries Society, 58(1):246–247, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moore:1929:AHW

[Moo29] Emmeline Moore. The aeration of hatchery water supplies and the incidence of disease. Transactions of the American Fisheries Society, 59(1):195–196, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Munro:1928:GWG

[Mun28] A. J. Munro. The glaucous-winged gull in its relation to fishing interests in British Columbia. Transactions of the American Fisheries Society, 58(1):161–164, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nevin:1921:FYE

[Nev21] James Nevin. Fifty years’ experience in fish culture. Transactions of the American Fisheries Society, 50(1):212–220, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nightingale:1927:PPS

[Nig27] H. W. Nightingale. Pollution problems in the State of Washington and their solution. Transactions of the American Fisheries Society, 57(1):294–300, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nightingale:1928:EWS

[NL28] H. W. Nightingale and V. Loosanoff. The effects of waste sulphite liquor on early stages of the Chinook salmon, and means of prevention by disposal methods. Transactions of the American Fisheries Society, 58(1):232–244, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Norgore:1922:SEF

[Nor22] Martin Norgore. Salmon eggs as food for salmon fry. Transactions of the American Fisheries Society, 51(1):82–86, January 1922. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nurnberger:1928:LPA

[Nur28] Patience Kidd Nurnberger. A list of the plant and animal food of some fishes of Jay Cooke Park. Transactions of the American Fisheries Society, 58(1):175–177, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nurnberger:1929:PAA

[Nur29] Patience Kidd Nurnberger. Plant and animal associations in a lake. Transactions of the American Fisheries Society, 59(1):174–177, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

OMalley:1928:ASA

[O’M28] Henry O’Malley. Alaska salmon — an achievement in conservation. Transactions of the American Fisheries Society, 58(1):39–43, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Orr:1920:RFM

[Orr20] Arthur Orr. The retail fish market: Some suggestions for equipping and conducting it. Transactions of the American Fisheries Society, 49(3):107–114, June 1920. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Osburn:1921:SPM

[Osb21a] Raymond C. Osburn. The scientist and the practical man in fisheries work. Transactions of the American Fisheries Society, 50(1):69–82, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Osburn:1921:SGF

[Osb21b] Raymond C. Osburn. A survey of game fish conditions in Ohio. Transactions of the American Fisheries Society, 50(1):353–363, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Osburn:1923:NAC

[Osb23] Raymond C. Osburn. Natural and artificial conditions detrimental to the black basses. Transactions of the American Fisheries Society, 53(1):27–49, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paulhus:1924:CPI

[Pau24] Mr J. A. Paulhus. Consumption, not production, is industry’s chief problem. Transactions of the American Fisheries Society, 54(1): 138–147, January 1924. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pollock:1928:IPS

[Pol28] Charles R. Pollock. International Pacific salmon investigation federation. Transactions of the American Fisheries Society, 58(1):68–70, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pope:1923:MMR

[Pop23] T. E. B. Pope. Museum methods in relation to the fisheries society. Transactions of the American Fisheries Society, 53(1):55–61, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Powers:1921:APU

[Pow21] Edwin B. Powers. Antagonism and its possible utility in polluted waters. Transactions of the American Fisheries Society, 50(1):293– 296, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prince:1920:WDS

[Pri20] Edward E. Prince. Why do salmon ascend from the sea? Transactions of the American Fisheries Society, 49(3):125–140, June 1920. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prince:1921:FYF

[Pri21] Edward E. Prince. Fifty years of fishery administration in Canada. Transactions of the American Fisheries Society, 50(1):163–186, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prince:1923:ICA

[Pri23] E. E. Prince. Irrigation canals as an aid to fisheries development in the West. Transactions of the American Fisheries Society, 52(1): 157–165, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prince:1924:WKE

[Pri24] E. E. Prince. What kind of education do fishermen need. Transactions of the American Fisheries Society, 54(1):25–28, January 1924. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Radcliffe:1921:NFA

[Rad21] Lewis Radcliffe. Notes on functions and activities of the division of fishery industries of the U.S. Bureau of Fisheries. Transactions of the American Fisheries Society, 50(1):106–119, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Radcliffe:1926:AWF

[Rad26] Lewis Radcliffe. Aquiculture–or water farming. Transactions of the American Fisheries Society, 56(1):83–90, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Radcliffe:1927:SA

[Rad27] Lewis Radcliffe. The status of aquiculture. Transactions of the American Fisheries Society, 57(1):171–176, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Radcliffe:1929:SGL

[Rad29] Lewis Radcliffe. Status of the Great Lakes fisheries. Transactions of the American Fisheries Society, 59(1):45–52, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reuling:1920:EAR

[Reu20] F. H. Reuling. Experiments in the artificial rearing of fresh-water mussels in troughs under conditions of control. Transactions of the American Fisheries Society, 49(3):153–155, June 1920. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robertson:1922:FPP

[Rob22a] Alexander Robertson. Further proof of the parent stream theory. Transactions of the American Fisheries Society, 51(1):87–90, January 1922. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robertson:1922:SOG

[Rob22b] Alexander Robertson. Some observation on the growth of young sockeyes. Transactions of the American Fisheries Society, 51(1):

91–96, January 1922. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rodd:1924:BLN

[Rod24a] J. A. Rodd. Barren lakes or natural nurseries for fry. Transactions of the American Fisheries Society, 54(1):161–162, January 1924. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rodd:1924:SDF

[Rod24b] J. A. Rodd. Sketch of the development of fish culture in Canada. Transactions of the American Fisheries Society, 54(1):148–160, January 1924. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scofield:1928:PSC

[Sco28] E. C. Scofield. Preliminary studies on the California striped bass. Transactions of the American Fisheries Society, 58(1):139–145, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1921:ASI

[Smi21a] Hugh M. Smith. The Alaska fur seal: An international asset. Transactions of the American Fisheries Society, 50(1):83–91, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1921:SRO

[Smi21b] Hugh M. Smith. Some recent observations on the freshwater EEL. Transactions of the American Fisheries Society, 50(1):297–300, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1926:SFW

[Smi26] Hugh M. Smith. Some fresh-water fishes of Siam. Transactions of the American Fisheries Society, 56(1):213–226, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snyder:1923:BBV

[Sny23] J. P. Snyder. Black bass versus netting of coarse fish. Transactions of the American Fisheries Society, 53(1):201–206, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snyder:1926:BBW

[Sny26] J. P. Snyder. Black bass — why? Transactions of the American Fisheries Society, 56(1):53–78, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1923:BSI

[Sur23] Thaddeus Surber. Biological surveys and investigations in Minnesota. Transactions of the American Fisheries Society, 52(1):225– 238, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1925:MUR

[Sur25] Thaddeus Surber. Methods used in restocking Minnesota lakes and streams, with comments on the danger attending over-stocking certain waters. Transactions of the American Fisheries Society, 55(1): 63–71, January 1925. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1929:USU

[Sur29a] Eugene W. Surber. The utilization of sloughs in the Upper Mississippi wild life and fish refuge as fish ponds. Transactions of the American Fisheries Society, 59(1):106–113, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1929:FCM

[Sur29b] Thaddeus Surber. Fish culture in Minnesota, past, present and future. Transactions of the American Fisheries Society, 59(1):224– 233, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Synnestvedt:1928:WWF

[Syn28] Harold Synnestvedt. The work of the Western Food and Game Fish Protective Association. Transactions of the American Fisheries Society, 58(1):78–79, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Taylor:1921:PIP

[Tay21] Harden F. Taylor. Principles involved in the preservation of fish by salt. Transactions of the American Fisheries Society, 50(1):120–152, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Titcomb:1928:NRG

[TCCM28] John W. Titcomb, Eben W. Cobb, Mary F. Crowell, and C. M. McCay. The nutritional requirements and growth rates of brook trout. Transactions of the American Fisheries Society, 58(1):205– 231, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Titcomb:1929:RVP

[TCCM29] John W. Titcomb, Eben W. Cobb, Mary F. Crowell, and C. M. Mc-Cay. The relative value of plant and animal by-products as feeds for brook trout and the basic nutritional requirements of brook trout in terms of proteins, carbohydrates, vitamins, inorganic elements, and roughage. Transactions of the American Fisheries Society, 59(1): 126–145, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:1925:SCB

[Tho25] W. T. Thompson. Scotty comes back. Transactions of the American Fisheries Society, 55(1):92–101, January 1925. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:1928:WIF

[Tho28] Will F. Thompson. The work of the International Fisheries Commission upon the Halibut. Transactions of the American Fisheries Society, 58(1):123–131, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:1929:SFC

[Tho29a] David H. Thompson. A survey of the fishes of Champaign County, Illinois. Transactions of the American Fisheries Society, 59(1):250– 255, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:1929:RL

[Tho29b] W. T. Thompson. A rival to liver. Transactions of the American Fisheries Society, 59(1):168–173, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tiffany:1921:GSR

[Tif21] L. H. Tiffany. The gizzard shad in relation to plants and game fishes. Transactions of the American Fisheries Society, 50(1):381– 386, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tilden:1929:PMD

[Til29] Josephine E. Tilden. “Plant material and debris”: The algal food of fishes. Transactions of the American Fisheries Society, 59(1): 178–187, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Titcomb:1921:SFC

[Tit21] John W. Titcomb. Some fish-cultural notes. Transactions of the American Fisheries Society, 50(1):200–211, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Titcomb:1922:GFL

[Tit22] John W. Titcomb. Growth of fish and location of hatcheries. Transactions of the American Fisheries Society, 51(1):65–69, January 1922. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Titcomb:1926:FRF

[Tit26] John W. Titcomb. Forests in relation to fresh water fishes. Transactions of the American Fisheries Society, 56(1):122–129, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Titcomb:1927:FWS

[Tit27] John W. Titcomb. The fresh-water shrimp for replenishing food in trout streams. Transactions of the American Fisheries Society, 57(1):150–161, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Turner:1921:FYL

[TK21] C. L. Turner and W. C. Kraatz. Food of young large-mouth black bass in some Ohio waters. Transactions of the American Fisheries Society, 50(1):372–380, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Toll:1925:ONP

[Tol25] Roger W. Toll. Our national parks. Transactions of the American Fisheries Society, 55(1):153, January 1925. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Travers:1921:FPO

[Tra21] John T. Travers. Fighting pollution in Ohio — a demonstration of methods. Transactions of the American Fisheries Society, 50(1): 279–292, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tulian:1920:LGP

[Tul20] E. A. Tulian. Louisiana–greatest in the production of shrimp — Penaeus setiferus. Transactions of the American Fisheries Society, 49 (3):115–124, June 1920. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tulian:1923:PSL

[Tul23] E. A. Tulian. The present status of the Louisiana shrimp industry. Transactions of the American Fisheries Society, 53(1):110–121, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tulian:1926:ISW

[Tul26] E. A. Tulian. Increase in the salt water shrimp catch from Louisiana waters. Transactions of the American Fisheries Society, 56(1):169– 174, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Taylor:1923:IPF

[TW23] Harden F. Taylor and Arthur W. Wells. Investigations in the preservation of fish nets and lines. Transactions of the American Fisheries Society, 52(1):35–73, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanAtta:1925:BRP

[Van25] C. H. Van Atta. Benefit of rearing ponds. Transactions of the American Fisheries Society, 55(1):72–79, January 1925. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanOosten:1929:SFP

[Van29] John Van Oosten. Some fisheries problems on the Great Lakes. Transactions of the American Fisheries Society, 59(1):63–85, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vickers:1923:OWM

[Vic23] H. W. Vickers. Oysters — the world’s most valuable seafood. Transactions of the American Fisheries Society, 52(1):201–205, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Viosca:1926:WPL

[Vio26] Percy Viosca, Jr. Water pollution in Louisiana. Transactions of the American Fisheries Society, 56(1):101–107, January 1926. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Viosca:1927:FCM

[Vio27] Percy Viosca, Jr. Flood control in the Mississippi Valley in its relation to Louisiana fisheries. Transactions of the American Fisheries Society, 57(1):49–64, January 1927. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Viosca:1928:FCL

[Vio28] Percy Viosca, Jr. Fish culture in Louisiana. Transactions of the American Fisheries Society, 58(1):165–166, January 1928. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Viosca:1929:FCL

[Vio29] Percy Viosca, Jr. Fish culture in Louisiana. Transactions of the American Fisheries Society, 59(1):207–216, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wallich:1925:RTF

[Wal25] Claudius Wallich. Rearing trout to the fingerling stage. Transactions of the American Fisheries Society, 55(1):129–134, January 1925. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ward:1921:APS

[War21a] Henry B. Ward. Atlantic and Pacific salmon. Transactions of the American Fisheries Society, 50(1):92–95, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ward:1921:SFM

[War21b] Henry B. Ward. Some features in the migration of the sockeye salmon and their practical significance. Transactions of the American Fisheries Society, 50(1):387–426, January 1921. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Webster:1929:PM

[Web29a] B. O. Webster. Propagation of muskellunge. Transactions of the American Fisheries Society, 59(1):202–203, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Webster:1929:ANP

[Web29b] B. O. Webster. To assist natural propagation in inland lakes. Transactions of the American Fisheries Society, 59(1):204–206, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wells:1921:APO

[Wel21] William Firth Wells. Artificial propagation of oysters. Transactions of the American Fisheries Society, 50(1):301–306, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Westerman:1929:PTP

[Wes29] F. A. Westerman. Progress in trout propagation in Michigan. Transactions of the American Fisheries Society, 59(1):234–237, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wheeler:1923:CF

[Whe23] Charles E. Wheeler. Commercial fisheries. Transactions of the American Fisheries Society, 52(1):206–209, January 1923. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wickliff:1921:FYS

[Wic21] Edward L. Wickliff. Food of young small-mouth black bass in Lake Erie. Transactions of the American Fisheries Society, 50(1):364– 371, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wiebe:1929:EVF

[Wie29] A. H. Wiebe. The effects of various fertilizers on plankton production. Transactions of the American Fisheries Society, 59(1):94–105, January 1929. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Williams:1920:WAF

[Wil20] J. Asakiah Williams. A word about Florida’s fish and her fisheries. Transactions of the American Fisheries Society, 49(4):159– 168, September 1920. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Willey:1921:EHC

[Wil21] Arthur Willey. The economic history of copepods. Transactions of the American Fisheries Society, 50(1):320–325, January 1921. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).